if a chemist wishes to male 3 liters of a 7% acid solution by mixing a 9% acid and a 4% acid solution . How many liters of each solution should the chemist use ?define variables write the equation that models the problem solve equation state the answer solution
Answer by Alan3354(69443)

Do it like this one. Only the numbers are different.
-----------
Amanda wants to make 8 gal. of a 20% saline solution by mixing together a 56% saline solution and a 8% saline solution. How much of each solution must she use?
================
e = amount of 8%
f = amount of 56%
---
e+ f = 8 (total solution)
8e + 56f = 20*8 (total saline)
---
e+ f = 8
e + 7f = 20
------------ Subtract
-6f = -12
f = 2
e = 6